Mr. Kim is a U.S. citizen who has been stationed at his employer’s Tokyo office for the last six years.

Required:

  1. Compute Mr. Kim’s AGI if his only income for the year was his $203,600 salary.
Answer #1

Since year has not been mentoned, Answer is given for 2020 and 2019 as well.

AGI = Salary - foreign earned income exclusion = 203600 - 107600 = 96000 based on year 2020 tax rates

AGI = Salary - foreign earned income exclusion = 203600 - 105900 = 97700 based on year 2019 tax rates